# Upper bounds of nodal sets for Gevrey regular parabolic equations

### v1

@@ Abstract.
We consider the size of the nodal set of the solution of the second order parabolic-type equation with Gevrey regular coefficients. We provide an upper bound as a function of time. The dependence agrees with a sharp upper bound when the coefficients are analytic. September 15, 2024
@@ 1. Introduction
in $\mathbb{T}^{d}\times(0,T)$, where $d\geq 1$, with given $v$ and $w$ and the initial data $u(\cdot,0)=u_{0}\in L^{2}(\mathbb{T}^{d})$. Under the assumption that $v$ and $w$ are $\frac{1}{\beta}$-Gevrey regular, we provide an upper bound on the $(d-1)$-dimensional Hausdorff measure of the nodal set. Our main result, Theorem 2.1, shows that for any $\beta\in(0,1]$ and $t\in(0,t_{0}]$, we have
@@
for some constant $C>0$ depending on the Dirichlet quotient of $u$, and the regularity of $v$ and $w$.
@@
In recent years, Logunov and Malinnikova achieved significant progress towards Yau’s conjecture for smooth manifolds [Lo1, Lo2, LoM1]. In [Lo1], Logunov proved Nadirashvili’s conjecture and showed the lower bound in Yau’s conjecture for compact $C^{\infty}$-smooth Riemannian manifold without boundary. In [LoM1], the authors established further estimates on the nodal sets of Laplace eigenfunctions; their result refined the upper bound in the Donnelly-Fefferman estimate when $d=2$ and yielded a lower bound when $d=3$. We refer the reader to the review of Yau’s conjecture in [LoM2] and for further results and references. Also, see [D, Do, HL, HS, L, LZ] for other results on the size of the nodal sets of solutions of partial differential equations.
@@
2. Gevrey regular parabolic problem and the main result
@@
where $\Omega=\mathbb{T}^{d}$ is the $d$-dimensional torus with side length $2\pi$ and $I=(0,t_{0}]$ is a given time interval. We require that $u$, $v$, and $w$ are periodic with respect to $\mathbb{T}^{d}$. We assume that $u_{0}$ is not identically zero. Suppose that there exist constants $M_{0},M_{1}\geq 1$ such that
@@
Suppose that $u(x,t)$ is the solution to (2.1) and (2.2) on $\mathbb{T}^{d}\times I$ with $u_{0}\in L^{2}(\mathbb{T}^{d})$ not identically zero. Assume that $v$ and $w$ satisfy (2.3)–(2.4) and (2.6)–(2.7). Then, for each $t\in I$, we have
@@
where $C>0$ is a constant depending on $q_{0}$, $M_{0}$, $M_{1}$, $K_{v}$ and $K_{w}$.
@@ 3. Proof of the main theorem
for all $t\in(0,t_{0}]$, where $P$ is a nonnegative polynomial.
@@
then $\delta$ is obtained by requiring
$\frac{1}{\mu^{2}}\geq\frac{C}{\mu_{0}^{2}}\log\frac{1}{\mu}+\frac{C(\beta(\mu)+1)}{\delta_{0}^{2}}$ (3.2)
@@
$\delta=\frac{\mu_{0}\sqrt{\frac{t}{2}}}{K\log^{\frac{1}{2}}\frac{1}{\mu_{0}}},$
@@
$\|u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}\leq C^{K^{2}\log(\frac{1}{\mu_{0}})\frac{1}{t}}\|u(t)\|_{L^{2}(B(\mu_{0}))}^{2},$ (3.4)
@@
$\begin{split}&\frac{1}{2}\frac{d}{dt}\|e^{\alpha tA^{\beta/2}}u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}+\|A^{\frac{1}{2}}e^{\alpha tA^{\beta/2}}u\|_{L^{2}(\mathbb{T}^{d})}^{2}\\ &=(e^{\alpha tA^{\beta/2}}(w\cdot\nabla u),e^{\alpha tA^{\beta/2}}u)+\alpha\|A^{\beta/4}e^{\alpha tA^{\beta/2}}u\|_{L^{2}}^{2}+(e^{\alpha tA^{\beta/2}}(vu),e^{\alpha tA^{\beta/2}}u)=I_{1}+I_{2}+I_{3}.\end{split}$ (3.5)
@@
Taking the inner product of (2.1) with $u$ and using $\|\nabla u\|^{2}_{L^{2}(\Omega)}=q_{D}(t)\|u\|_{L^{2}(\Omega)}^{2}$, where $q_{D}(t)$ is defined in (2.5), we obtain
[…]
@@
$\begin{split}\frac{1}{2}\frac{d}{dt}\|u\|^{2}_{L^{2}(\mathbb{T}^{d})}&\geq-(\|\mathop{\rm div}\nolimits w\|_{L^{\infty}(\mathbb{T}^{d})}+\|v(t)\|_{L^{\infty}(\mathbb{T}^{d})}+q_{D}(t))\|u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}\\ &\geq-(M_{1}+M_{0}+q_{0})\|u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}.\end{split}$
@@
Let $r>0$. Suppose that $N\subset B_{2r}$ is a subset of a countable union of $(d-1)$-dimensional $C^{1}$ graphs. Let $T=\{p_{\pm 1},\ldots,p_{\pm d}\}$, where $p_{j}\in B_{r/10d}(re_{j})$ with the standard basis $e_{j}$ in $\mathbb{R}^{d}$, for $j=\pm 1,\ldots,\pm d$. If $N$ has at most $n$ intersections with any line through any $p\in T$, then
@@
for some constant $C>0$. Fix $q\in N$. Note that there are at most $n$ unit normal vectors to $N$ at $q$. Let $\nu_{q}=(\mu_{1},\ldots,\mu_{d})$ be a unit normal vector to $N$ at $q$ and $|\mu_{k}|=\max\{|\mu_{1}|,\ldots,|\mu_{d}|\}$ for some $k\in\{1,2,\ldots,d\}$. We claim that there exists some $p\in T$ such that
@@
For each $p\in T$, denote by $F_{p}$ the collection of points $q\in N$ such that there exists a unit normal vector $\nu_{q}$ to $N$ at $q$ with (3.14) and (3.15). Since $N$ has at most $n$ intersections with any line through $p$, we infer that $F_{p}$ is comparable to $B_{4}(p)$ up to a universal constant depending on $d$. Namely, we have
$\mathcal{H}^{d-1}(F_{p})\leq Cn\mathcal{H}^{d-1}(B_{4}(p))\leq Cn,$
@@
Note that the minimal number of points needed in $T$ is $d+1$. Indeed, we use a simple fact that if the determinant of $d$ vectors is nonzero, then these $d$ vectors are non-coplanar. Consequently, the angle between $q-p$ and $\nu_{q}$ in (3.14) is bounded away from zero.
@@
Fix $j\in\{\pm 1,\ldots,\pm d\}$ and a line $l$ through $q_{j}$. We assume that $u$ has $n$ zeros in the line segment $l\cap B_{2r}(p)$, counting the multiplicity. We rotate the coordinate so that $u$ is a function of a single variable and we use $u^{(m)}$ to denote the $m$-th derivative of $u$. Let $\{x_{1},\ldots,x_{k}\}$ be $k$ distinct zeros in $l\cap B_{2r}(p)$ with multiplicities $m_{1},\ldots,m_{k}\geq 1$ and $m_{1}+\cdots+m_{k}=n$. By the Hermite interpolation theorem, there exists a unique interpolation polynomial $P_{n-1}$ of $u$, of degree less than or equal to $n-1$, satisfying
@@
Let $n_{0}$ be the largest integer less than or equal to $12K^{2}\log(\frac{1}{r})\frac{1}{t}+2d+$ 1. In the left side of (3.22) we replace $n$ with $n_{0}$, and the resulting quantity is bounded from above by
@@
where $M\geq 1$ is a sufficiently large constant depending on $C_{0},C_{1},C_{2}$, $C_{3}$ and $K$. Then the quantity in (3.23) is less than 00 for all $t\in(0,t_{0}]$, since $\beta/(2-\beta)\leq 1$. Consequently, the value $n=n_{0}$ does not satisfy (3.22) and thus does not satisfy (3.20). We infer that $u$ has less than $n_{0}$ zeros in $l\cap B_{2r}(p)$ for fixed $r>0$ as in (3.24).
For $t\in(0,t_{0}]$, since the order of vanishing for $u(x,t)$ is finite for any $x\in\mathbb{T}^{d}$, we infer from [M, Claim 2] that $N(t)$ is contained in a union of $(d-1)$-dimensional $C^{1}$ graphs. Without loss of generality, we may assume that the union is taken on a countable set, as otherwise we consider any countable subset of the uncountable union and take the supremum. From Lemma 3.4 it follows that

### v2

@@ Abstract.
We consider the size of the nodal set of the solution of the second order parabolic-type equation with Gevrey regular coefficients. We provide an upper bound as a function of time. The dependence agrees with a sharp upper bound when the coefficients are analytic.
@@ 1. Introduction
in $\mathbb{T}^{d}\times(0,T)$, where $d\geq 1$, with given $v$ and $w$ and the initial data $u(\cdot,0)=u_{0}\in H^{1}(\mathbb{T}^{d})$. Under the assumption that $v$ and $w$ are $\frac{1}{\beta}$-Gevrey regular, we provide an upper bound on the (d1)(d-1)-dimensional Hausdorff measure of the nodal set. Our main result, Theorem 2.1, shows that for any $\beta\in(0,1]$ and $t\in(0,t_{0}]$, we have
@@
for some constant $C>0$ depending on the initial condition of $u_{0}$, and the regularity of $v$ and $w$, assuming $t_{0}\in(0,1/2]$.
@@
In recent years, Logunov and Malinnikova achieved significant progress towards Yau’s conjecture for smooth manifolds [Lo1, Lo2, LoM1]. In [Lo2], Logunov proved Nadirashvili’s conjecture and showed the lower bound in Yau’s conjecture for compact $C^{\infty}$-smooth Riemannian manifold without boundary. In [LoM1], the authors established further estimates on the nodal sets of Laplace eigenfunctions; their result refined the upper bound in the Donnelly-Fefferman estimate when $d=2$ and yielded a lower bound when $d=3$. We refer the reader to the review of Yau’s conjecture in [LoM2] and for further results and references. Also, see [D, Do, HL, HJ, HS, KL, L, LZ, TY] for other results on the size of the nodal sets of solutions of partial differential equations.
@@
2. G·evrey regular parabolic problem and the main result
@@
where $\Omega=\mathbb{T}^{d}$ is the $d$-dimensional torus with side length $2\pi$ and $I=(0,t_{0}]$ is a given time interval, where $t_{0}\in(0,1/2]$. We require that $u$, $v$, and $w$ are $2\pi$-periodic. We assume that $u_{0}\in H^{1}(\mathbb{T}^{d})$ is not identically zero. Suppose that there exist constants $M_{0},M_{1}\geq 1$ such that
@@
Suppose that $u(x,t)$ is the solution to (2.1) and (2.2) on $\mathbb{T}^{d}\times I$ with $u_{0}\in H^{1}(\mathbb{T}^{d})$ not identically zero. Assume that $v$ and $w$ satisfy (2.3)–(2.4) and (2.7)–(2.8). Then, for each $t\in I$, we have
@@
where $C>0$ is a constant depending on $M_{0}$, $M_{1}$, $K_{v}$ and $K_{w}$.
@@
Here we comment on the boundedness of the Dirichlet quotient $q_{\text{D}}$ in (2.5). As in [FS, p. 463], we have
$\frac{1}{2}\frac{d}{dt}q_{\text{D}}+\left\|(-\Delta-q_{\text{D}})\frac{u}{\|u\|_{L^{2}}}\right\|^{2}=\frac{1}{\|u\|^{2}}\int_{\Omega}(w\cdot\nabla u+vu)(-\Delta-q_{\text{D}})u$
which implies by Cauchy-Schwarz inequality
$\frac{d}{dt}q_{\text{D}}\leq C(\|w\|_{L^{\infty}}^{2}q_{\text{D}}+\|v\|_{L^{\infty}}^{2}),$
and this gives an explicit bound on $q_{\text{D}}(t)$ for $t\in(0,t_{0}]$ depending on $q_{\text{D}}(t_{0})$.
@@ 3. Proof of the main theorem
for all $t\in(0,t_{0}]$, where $P$ is a nonnegative polynomial and $q_{0}$ is as in (2.6).
@@
then $\mu$ is obtained by requiring
$\frac{1}{\mu^{2}}\geq\frac{C}{\mu_{0}^{2}}\log\frac{1}{\mu}+\frac{C(\beta(\mu)+1)}{\mu_{0}^{2}}$ (3.2)
@@
$\mu=\frac{\mu_{0}\sqrt{\frac{t}{2}}}{K\log^{\frac{1}{2}}\frac{1}{\mu_{0}}},$
@@
$\|u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}\leq C^{K^{2}\log(\frac{1}{\mu_{0}})\frac{1}{t}}\|u(t)\|_{L^{2}(B_{\mu_{0}})}^{2},$ (3.4)
@@
$\begin{split}&\frac{1}{2}\frac{d}{dt}\|e^{\alpha tA^{\beta/2}}u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}+\|A^{\frac{1}{2}}e^{\alpha tA^{\beta/2}}u\|_{L^{2}(\mathbb{T}^{d})}^{2}\\ &=(e^{\alpha tA^{\beta/2}}(w\cdot\nabla u),e^{\alpha tA^{\beta/2}}u)+\alpha\|A^{\beta/4}e^{\alpha tA^{\beta/2}}u\|_{L^{2}}^{2}+(e^{\alpha tA^{\beta/2}}(vu),e^{\alpha tA^{\beta/2}}u)\\ &=I_{1}+I_{2}+I_{3}.\end{split}$ (3.5)
@@
Taking the inner product of (2.1) with $u$ and using $\|\nabla u\|^{2}_{L^{2}(\Omega)}=q_{\text{D}}(t)\|u\|_{L^{2}(\Omega)}^{2}$, where $q_{\text{D}}(t)$ is defined in (2.5), we obtain
$\frac{1}{2}\frac{d}{dt}\|u\|^{2}_{L^{2}(\mathbb{T}^{d})}+q_{\text{D}}(t)\|u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}=-\frac{1}{2}\int_{\mathbb{T}^{d}}(\mathop{\rm div}\nolimits w)u^{2}\,dx+\int_{\mathbb{T}^{d}}vu^{2}\,dx.$
@@
$\begin{split}\frac{1}{2}\frac{d}{dt}\|u\|^{2}_{L^{2}(\mathbb{T}^{d})}&\geq-(\|\mathop{\rm div}\nolimits w\|_{L^{\infty}(\mathbb{T}^{d})}+\|v(t)\|_{L^{\infty}(\mathbb{T}^{d})}+q_{\text{D}}(t))\|u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}\\ &\geq-(M_{1}+M_{0}+q_{0})\|u(t)\|_{L^{2}(\mathbb{T}^{d})}^{2}.\end{split}$
@@
Let $r>0$. Suppose that $N\subset B_{2r}$ is a subset of a countable union of (d1)(d-1)-dimensional $C^{1}$ graphs. Let $T=\{p_{\pm 1},\ldots,p_{\pm d}\}$, where $p_{j}\in B_{r/10d}(re_{j})$ with the standard basis $e_{j}$ in $\mathbb{R}^{d}$, for $j=\pm 1,\ldots,\pm d$. If $N$ has at most $n$ intersections with any line through any $p\in T$, then
@@
for some constant $C>0$. Fix $q\in N$. We claim that there are at most $n$ pairwise non-parallel unit normal vectors to $N$ at $q$. Suppose, for the sake of contradiction, that there are $n+1$ pairwise non-parallel unit normal vectors at $q$. There exists $\kappa>0$ and a line $l$ through a point $p\in T$ such that
$\mathop{\rm card}\nolimits(N\cap B_{\kappa}(q)\cap l)\geq n+1.$
The existence of such a line is obtained by perturbing the line connecting $p$ and $q$ and applying the implicit function theorem. This contradicts the assumption that every line through $p\in T$ intersects $N$ in at most $n$ points. Let $\nu_{q}=(\mu_{1},\ldots,\mu_{d})$ be a unit normal vector to $N$ at $q$ and $|\mu_{k}|=\max\{|\mu_{1}|,\ldots,|\mu_{d}|\}$ for some $k\in\{1,2,\ldots,d\}$. We claim that there exists some $p\in T$ such that
@@
For each $p\in T$, denote by $F_{p}$ the collection of points $q\in N$ such that there exists a unit normal vector $\nu_{q}$ to $N$ at $q$ with (3.14) and (3.15). Since $N$ has at most $n$ intersections with any line through $p$, we infer that the (d1)(d-1)-dimensional Hausdorff measure of $F_{p}$ is bounded by $n$ times the (d1)(d-1)-dimensional Hausdorff of the sphere $\mathbb{S}^{d-1}_{4}(p)=\{x\in\mathbb{R}^{d}:|x-p|=4\}$, up to a universal constant depending on $d$. Namely, we have
$\mathcal{H}^{d-1}(F_{p})\leq Cn\mathcal{H}^{d-1}(\mathbb{S}^{d-1}_{4}(p))\leq Cn,$
@@
Note that the minimal number of points needed in $T$ is $d+1$. Indeed, we use a simple fact that if the determinant of $d$ vectors is nonzero, then these $d$ vectors are non-coplanar. Consequently, the angle between qpq-p and $\nu_{q}$ in (3.14) is bounded away from zero.
@@
Fix $j\in\{\pm 1,\ldots,\pm d\}$ and a line $l$ through $q_{j}$. We assume that $u$ has $n$ zeros in the line segment $l\cap B_{2r}(p)$, counting the multiplicity. We rotate the coordinate so that $u$ is a function of a single variable and we use $u^{(m)}$ to denote the $m$-th derivative of $u$. Let $\{x_{1},\ldots,x_{k}\}$ be $k$ distinct zeros in $l\cap B_{2r}(p)$ with multiplicities $m_{1},\ldots,m_{k}\geq 1$ and $m_{1}+\cdots+m_{k}=n$. By the Hermite interpolation theorem, there exists a unique interpolation polynomial $P_{n-1}$ of $u$, of degree less than or equal to n1n-1, satisfying
@@
Let $n_{0}$ be the largest integer less than or equal to $2K^{2}\log(\frac{1}{r})\frac{1}{t}+2d+1$. In the left side of (3.22) we replace $n$ with $n_{0}$, and the resulting quantity is bounded from above by
@@
where $M\geq 1$ is a sufficiently large constant depending on $C_{0},C_{1},C_{2}$, $C_{3}$ and $K$. Then the quantity in (3.23) is less than 0 for all $t\in(0,t_{0}]$, since $\beta/(2-\beta)\leq 1$. Consequently, the value $n=n_{0}$ does not satisfy (3.22) and thus does not satisfy (3.20). We infer that $u$ has less than $n_{0}$ zeros in $l\cap B_{2r}(p)$ for fixed $r>0$ as in (3.24). Assume, for the sake of contradiction, that the order of vanishing of $u(x,t)$ is infinite for some $x\in\mathbb{T}^{d}$ and some $t\in(0,t_{0}]$. Then, for the line $l$ passing $q_{j}$ and $x$, the restriction of $u$ to $l\cap B_{2r}(x)$ has a zero at $x$ of infinite multiplicity. This contradicts with the fact that the multiplicity of zeros of $u$ is bounded by $n_{0}$ in $l\cap B_{2r}(p)$, uniformly for any $p\in\mathbb{T}^{d}$.
For $t\in(0,t_{0}]$, since the order of vanishing for $u(x,t)$ is finite for any $x\in\mathbb{T}^{d}$, we infer from [M, Claim 2] that $N(t)$ is contained in a union of (d1)(d-1)-dimensional $C^{1}$ graphs. Without loss of generality, we may assume that the union is taken on a countable set, as otherwise we consider any countable subset of the uncountable union and take the supremum. From Lemma 3.4 it follows that